Make y the subject of the equation below
y+9
___ = 2
4

Found 2 solutions by nanakwame, edjones:
Answer by nanakwame(5)

(y 9)/4=2
y 9=2กม4
y 9=8
y=8-9
y=-1

Answer by edjones(8007)

(y+9)/4=2
y+9=8 multiply each side of the equation by 4.
y=-1 add -9 to each side.
.
Ed